# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

### v4

@@ A.5 Clustering Results
Tables A13 and A14 contain the manually annotated Cluster Theme, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A11 and A12.
@@
Table A11: Themes ($n=26$) and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood. The cluster Lack of shame was removed since the reason behind shame was not evident. Clusters Lack of accountability and Inappropriate social behavior were merged together (and named ‘‘Lack of accountability’’) due to overlapping normative expectations.
@@
Figure A1: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female-Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates significant difference i.e., $p>0.05$

### v5

@@ A.5 Clustering Results
Tables A13 and A14 contain the manually annotated Cluster Themes, the total number of samples in each cluster, and Bollywood vs Hollywood distribution. The distance was set to $5$ and the duplicates were removed. The theme and top three examples demonstrating its meaning are provided in Tables A11 and A12.
@@
Table A11: Themes ($n=26$) and top-3 examples for clusters obtained after agglomerative clustering of shame-related dialogues from Bollywood and Hollywood. The cluster Lack of shame was removed since the reason behind shame was not evident. Clusters Lack of accountability and Inappropriate social behavior were merged together (and named “Lack of accountability”) due to overlapping normative expectations.
@@
Figure A1: Gender-wise differences in themes associated with normative expectations from pride-related discourse in Bollywood and Hollywood. $\Delta(Female-Male)$ is the difference between normalized dialogues attributed to females under a theme and to males under the same theme (as in eq. 1 but for dimension gender). * indicates significant difference i.e., $p<0.05$